A flywheel rotates with a constant retardation due to breaking from t=0to t=10 seconds. it made 400 revolutions . At time t=8 sec , it's angular velocity was 45πrad/sec. Determine

1 . Value of constant ratardation.

2. Total time taken to come to rest.

3. Total revolution made till it comes to rest.


1
Expert's answer
2021-05-07T08:12:27-0400

"t_o = 0s\\
\omega_o= ?\\
t_1 = 8s \\
\omega_1 = 40π rad/s\\
t_2 = 10s\\
\omega_2 = ?"


Retardation is constant throughout,

"\therefore \alpha = \dfrac{\omega_1- \omega_o}{t_1- t_o}=\dfrac{\omega_2- \omega _1}{t_2- t_1}"


"\dfrac{40π-\omega_o}{8} = \dfrac{\omega_2 - 40π}{2}"


"\dfrac{40π-\omega_o}{4} = {\omega_2 - 40π}"


"40π-\omega_o= 4{\omega_2 - 160π}\\
4\omega_2+ \omega_o=200π---(i)"



"\theta =\dfrac{\omega_o +\omega_2}2× t\\"


"400(2π) =\dfrac{\omega_o +\omega_2}2× 10"


"800π = \dfrac{\omega_o +\omega_2}2× 10"


"160π= \omega_o +\omega_2---(ii)"


Comparing equation (i) with equation (ii),

"\omega_o = 461\ rad/s, \quad \omega_2= 52\ rad/s"


"\theta= \omega_o t + 0.5\alpha t²\\
800π =461(10) + 0.5\alpha(10)^ 2"


"\alpha = -42\ rad/s²"